Starting with the graph of f(x) = 2^x, write the equation of the graph that results from reflecting f(x) about the y-axis. y = ?
Answer #1
f(x) =2^x,,,,,on reflection about y axis,,f(x) =2^(-x) . { logic,,,f(x) =y,,,, reflection about y axis,,,has same y value ,,but . negative x value . plot graph to check .
answered by: Coroline